Evaluate the integral of (dt)/(cos² t)


1
Expert's answer
2021-03-09T03:31:54-0500

"\int\frac{dt}{\cos^2{\ \left(t\right)}}"

 "let\ x=\tan{\left(t\right)} → \cos{\left(t\right)}=\frac{1}{\sqrt{1+x^2}},\sin{\left(t\right)}=\frac{x}{\sqrt{1+x^2}}\"

"dx=\sec^2{\left(t\right)\ dt}"

"dt=\cos^2{\left(t\right)\ dx\ }"

"dt=\frac{dx}{1+x^2}"

"\int\frac{dt}{\cos^2{\left(t\right)}}=\int\frac{1+x^2}{1}\frac{dx}{1+x^2}\ =\ \int dx"

"x+c"

"\tan{\left(t\right)}+c"